Define a binary relation P from R to R as follows: for all real numbers x and y, (x,y)∈P⇔x=y^2. Is P a function? Explain.


1
Expert's answer
2021-06-02T11:05:23-0400

Taking into account that "4=2^2" and "4=(-2)^2", we conclude that "(4,2)\in P" and "(4,-2)\in P". Therefore, the element "x=4\in \mathbb R" corresponds to two elements "y=2\in\mathbb R" and "y=-2\in\mathbb R", and hence "P:\mathbb R\to\mathbb R" is not a function "y=f(x)."


On the other hand, for each element "y\in\mathbb R" there exists a unique element "x=y^2\in\mathbb R". Consequently, "P:\mathbb R\to \mathbb R, P(y)=y^2," is a function of variable "y."